OpenStudy (anonymous):

simplify sqrt 20b^3c^4

OpenStudy (anonymous):

\[\sqrt{20b^3c^4}=\sqrt{4b^2c^4}*\sqrt{5b}=2bc^2\sqrt{5b}\]